OpenStudy (anonymous):

Pleasee help me with this math problem!! Write the equation of the circle with the given center an radius. Center: (0, 2) Radius: (square root of 13)

OpenStudy (anonymous):

@catalinaariza Do you know any circle equation?

OpenStudy (anonymous):

Well I have 4 options a. (x – 2)2 + y2 = 169 b. x2 + (y + 2)2 = 13 c. (x + 2)2 + y2 = 169 d. x2 + (y – 2)2 = 13

OpenStudy (anonymous):

(x-h)^2 + (y-k)^2 = r^2 Thats the equation of a circle so plug in your numbers

OpenStudy (anonymous):

@catalinaariza All you do is PLUG the numbers in, to find the right result yourself!

OpenStudy (anonymous):

Okay thankssssss :)